# Live Video Captioning

### v1

@@ IV-E Qualitative results
IV-F Demo
In this section, we showcase the system operating in real-time and provide information about its performance in this regard. We have implemented a demo that allows for the direct processing of a video stream from a camera using our LVC model.
To achieve processing speed close to real-time and provide captions with good quality, we have had to make the following interventions in the LVC model. The demonstrator works by directly accessing a video stream from a webcam. Once launched, the model will begin generating captions on the video stream immediately and continuously, displaying them on the screen. We have implemented a multiprocessing solution with two threads running simultaneously. The multiprocessing implementation is essential to ensure that while a caption is being generated, frames continue to be captured to avoid losing information. The first thread is responsible for capturing frames and displaying the images and captions on the screen. The second thread processes sets of frames to produce the captions using our LVC model. This caption generation thread can be parameterized to define both the parameter $\Delta t$, i.e., the length of the minimum video segment to be analyzed, and a memory parameter $M$ that is maximum number of segments with length $\Delta t$ that we will keep in memory to produce captions in the demonstrator. With this memory, we ensure that the dense caption prediction LVC module can access more context, a larger portion of the video, without losing its essence as a live system, and generate more accurate captions. Figure 14 provides a graphical description of the implementation made for the demonstrator.
@@ IV-F Demo
As for processing speed, in Table VI, we report the average frames per second that our implementation is capable of processing and the average time it takes to generate a caption prediction by our LVC implementation. The camera we are working with provides a frame rate of 30 frames per second, so we can calculate the optimal length of the video segments to be used by the LVC system. We define $l$ as the length in frames of the input video clip: $l=2.38\cdot 30\approx 71$ If the video segments we introduce to the model have a duration of 71 frames, we will ensure that when the processing of one clip ends, the next one is introduced into the LCV model. This way, no frames are discarded. All these tests for the demo were performed on a laptop running Ubuntu 18.04 operating system, equipped with an Intel Core i7 processor, and an integrated NVIDIA Quadro RTX 5000 MaxQ graphics card. Note that higher speed can be achieved if a more powerful GPU is used.
@@
TABLE VI: Number of frames per second (FPS) and average prediction time reported by our LVC demo system.
V Conclusions
@@
Live video captioning is a novel and challenging problem that has not been deeply investigated in the scientific literature. As we have shown, generating dense captions for live video streams is a much harder problem than one might conclude from results reported in previous works under more constrained settings, e.g., offline dense video captioning models. In fact, traditional evaluation metrics need to be updated to novel online versions that allow us to judge the actual live performance of the LVC models. In this work, we have formalized the problem of LVC for the first time, proposed new metrics tailored to it, and introduced an LVC model capable of integrating transformer-based attention mechanisms with a caption filtering module for video streams received as input. Results of our model as well as an evaluation kit with the novel metrics integrated are made publicly available to encourage further research on LVC on realistic data: https://github.com/gramuah/lvc. We hope to encourage more researchers to look into the challenging yet very practical task of LVC. This work not only advances the understanding of LVC but also opens new avenues for real-time video understanding and accessibility applications in dynamic environments.
Acknowledgments
@@
This research was partially funded by projects: NAVISOCIAL, with reference 2023/00405/001 from the University of Alcalá; AIRPLANE, with reference PID2019-104323RB-C3 from the Ministry of Science and Innovation of Spain. We appreciate the assistance of Walfrido González Molina in generating some figures.

### v2

@@ 4.5 Qualitative analysis
4.6 Demo
In this section, we showcase the system in operation with a live video stream and highlight its performance. We have implemented a demo that allows for the direct processing of a video stream from a camera using our LVC model.
To achieve processing speed and provide captions with good quality, we have had to make the following interventions in the LVC model. The demonstrator works by directly accessing a video stream from a webcam. Once launched, the model will begin generating captions on the video stream immediately and continuously, displaying them on the screen. We have implemented a multiprocessing solution with two threads running simultaneously. The multiprocessing implementation is essential to ensure that while a caption is being generated, frames continue to be captured to avoid losing information. The first thread is responsible for capturing frames and displaying the images and captions on the screen. The second thread processes sets of frames to produce the captions using our LVC model. This caption generation thread can be parameterized to define both the parameter $\Delta t$, i.e., the length of the minimum video segment to be analyzed, and a memory parameter $M$ that is maximum number of segments with length $\Delta t$ that we will keep in memory to produce captions in the demonstrator. With this memory, we ensure that the dense caption prediction LVC module can access more context, a larger portion of the video, without losing its essence as a live system, and generate more accurate captions. Figure 15 provides a graphical description of the implementation made for the demonstrator.
@@ 4.6 Demo
As for processing speed, in Table 7, we report the average frames per second that our implementation is capable of processing and the average time it takes to generate a caption prediction by our LVC implementation. The camera we are working with provides a frame rate of 30 frames per second, so we can calculate the optimal length of the video segments to be used by the LVC system. We define $l$ as the length in frames of the input video clip: $l=2.38\cdot 30\approx 71$ If the video segments we introduce to the model have a duration of 71 frames, we will ensure that when the processing of one clip ends, the next one is introduced into the LCV model. This way, no frames are discarded. All these tests for the demo were performed on a laptop running Ubuntu 18.04 operating system, equipped with an Intel Core i7 processor, and an integrated NVIDIA Quadro RTX 5000 MaxQ graphics card. Note that higher speed can be achieved if a more powerful GPU is used.
@@
Table 7: Number of frames per second (FPS) and average prediction time reported by our LVC demo system.
For the LVC problem, it is crucial to assess the performance of our model on devices with limited computational capacity, such as those used in embedded systems, robotics, IoT, or Edge Computing. To this end, we evaluated our LVC model on an NVIDIA Jetson Xavier-NX board, a platform equipped with 384-core NVIDIA Volta GPUs and 48 Tensor Cores, capable of operating with a power consumption as low as 20W. This setup allowed us to analyze the model’s suitability for resource-constrained environments. To conduct these performance tests, we adapted the developed demo to operate efficiently on the NVIDIA Jetson Xavier-NX. First, we compiled the necessary libraries specifically for optimal use of the ARM64 architecture and the integrated GPU of the Jetson Xavier-NX board. After embedding the application into the platform, we analyzed the system’s performance, adjusting the parameter $\Delta t$ as shown in Table 8. One key observation is that our implementation achieves a frame rate exceeding 0.8 FPS, which could be sufficient for many applications requiring an embedded LVC system. Notably, the maximum $\Delta t$ value that the Jetson platform can process is 70 frames. It is worth highlighting that the processing of TSP features represents the main bottleneck. In percentage terms, the time spent generating these features consistently exceeds the time required for generating dense captions and filtering when $\Delta t>16$ frames.
$\Delta t$ FPS Time of features (sec.) Time of LVC (sec.)
16 0.87 0.39 0.75
48 0.51 1.15 0.75
70 0.11 7.6 2.4
Table 8: Performance evaluation of our LVC model on an NVIDIA Jetson Xavier-NX board. We report: number of frames per second (FPS), and average time for the compuation of features and captions.
5 Conclusions
@@
Live video captioning is a novel and challenging problem that has not been deeply investigated in the scientific literature. As we have shown, generating dense captions for live video streams is a much harder problem than one might conclude from results reported in previous works under more constrained settings, e.g., offline dense video captioning models. In fact, traditional evaluation metrics need to be updated to novel online versions that allow us to judge the actual live performance of the LVC models. In this work, we have formalized the problem of LVC for the first time, proposed new metrics tailored to it. We introduced an LVC model capable of integrating transformer-based attention mechanisms with a caption filtering module for video streams received as input. Results of our model as well as an evaluation kit with the novel metrics integrated are made publicly available to encourage further research on LVC on realistic data: https://github.com/gramuah/lvc. We hope to encourage more researchers to look into the challenging yet very practical task of LVC. This work enhances our understanding of LVC and paves the way for live video understanding and accessibility applications in dynamic environments.
As future work, we consider the following options. The first would be to allow our LVC model to improve the predictions of the captions, using a memory mechanism, where the model takes into account past predictions. Technically, for example, recurrent networks could be incorporated into the model. A second line of improvement would consist of incorporating explainability techniques to improve the interpretability of our LVC model. Inspired by advancements in Explainable Artificial Intelligence (XAI) [44], such as attention visualization and feature importance analysis, we aim to provide insights into the model’s decision-making process. These methods could help bridge the gap between black-box models and transparent systems, enhancing trust and usability in applications requiring streaming video analysis. Finally, the future capabilities LVC systems are closely tied to advancements in sensor technology and video analysis techniques. Improvements in camera sensors, such as higher frame rates, and low-light performance, can provide richer data for more accurate caption generation. Additionally, emerging video analysis techniques, including multimodal feature fusion and lightweight neural architectures, may enhance the efficiency and robustness of LVC systems, particularly for real-time applications. While these developments are beyond the scope of this study, they represent promising directions for expanding the applicability and effectiveness of LVC models in diverse scenarios, including robotics, surveillance, and assistive technologies.
Declarations
  • Acknowledgements: We appreciate the assistance of Walfrido González Molina in generating some figures.
@@
  • Funding: This research was partially funded by projects: NAVISOCIAL, with reference 2023/00405/001 from the University of Alcalá; NAVIGATOR-D, with reference PID2023-148310OB-I00 from the Ministry of Science and Innovation of Spain.
  • Conflict of interest/Competing interests: The authors certify that they have no conflict of interest.
  • Ethics approval and consent to participate: Not applicable.
  • Consent for publication: Not applicable.
  • Data availability: The dataset used for all the experiments in this paper is ActivityNet Captions [13], which can be accessed at https://cs.stanford.edu/people/ranjaykrishna/densevid/.
  • Materials availability: Not applicable.
  • Code availability: Code to reproduce our results can be found at https://github.com/gramuah/lvc.
  • Author contribution: Conceptualization, all; methodology, R.J.L.-S. and E. B.-F.; software, R.J.L.-S., E. B.-F., N.N. and C.G.-Á.; validation, all; ; writing—original draft preparation, all; supervision, R.J.L.-S. and S. M.-B.; project administration, R.J.L.-S. and S. M.-B.; funding acquisition, R.J.L.-S. and S. M.-B..